# PRESERVE: Prefetching Model Weights and KV-Cache in Distributed LLM Serving

### v1

@@
Abstract
Large language models (LLMs) are widely used across various applications, but their substantial computational requirements pose significant challenges, particularly in terms of HBM bandwidth bottlenecks and inter-device communication overhead. In this paper, we present PRESERVE, a novel prefetching framework designed to optimize LLM inference by overlapping memory reads for model weights and KV-cache with collective communication operations. Through extensive experiments conducted on commercial AI accelerators, we demonstrate up to 1.6× end-to-end speedup on state-of-the-art, open-source LLMs. Additionally, we perform a design space exploration that identifies the optimal hardware configuration for the proposed method, showing a further 1.25× improvement in performance per cost by selecting the optimal L2 cache size. Our results show that PRESERVE has the potential to mitigate the memory bottlenecks and communication overheads, offering a solution to improve the performance and scalability of the LLM inference systems.
1 Introduction
Large language models (LLM) have been widely deployed across various services and application domains such as chat assistants [28], code generation [5], and knowledge retrieval [20]. While LLMs have demonstrated outstanding capabilities in these domains, their substantial computational requirements result in slow and costly inference. To address these challenges, specialized accelerators have been developed and deployed for LLM inference. Consequently, the performance, efficiency, and scalability of LLM inference accelerators are critical to the speed and cost of the LLM inference.
@@ 1 Introduction
In this paper, we address both of these two major challenges at once. We propose prefetching the model weights and KV-cache from off-chip memory to on-chip cache during the communication operations. This approach allows the communication operations to be overlapped with the memory read operations for the model weights and KV-cache, hiding the latency of the latter. Once the communication operations are completed, compute cores access the required data from the on-chip cache, no longer limited by off-chip memory bandwidth. As a result, the overall decoding phase of LLM inference is executed faster and more efficiently, reducing end-to-end latency and execution cost.
Unfortunately, current hardware schedulers and modern ML compilers cannot facilitate the proposed memory-communication overlap due to data dependencies between compute and communication operations in LLM inference. To address this issue, other prior studies have proposed fusing compute and communication operations[11, 39]. However, the effectiveness of these methods is limited as they can only fuse two consecutive operations, which is often insufficient to hide the entire communication latency. As a result, existing methods fail to provide an effective and comprehensive solution.
Although compute and communication cannot normally be overlapped due to data dependencies on input data, the model weights and KV-cache do not depend on the results of previous collective communication operations. Therefore, the memory read operations for the model weights and KV-cache can be overlapped without violating any data dependency between operations. To that end, we propose PRESERVE, a novel method that prefetches model weights and KV-cache from off-chip memory to L2 cache in parallel to the collective communication operations. To maximize the performance gains without the need for any modification in the user code, we implemented PRESERVE as a graph optimization framework. Our graph optimization scheme inserts prefetch commands into a given computation graph, using parallel streams to the collective communication operations. PRESERVE also keeps track of the prefetched data and estimates L2 cache usage during compilation time in a way that the total prefetched data size does not exceed a predefined upper limit to prevent cache pollution.
@@ 2 Background
Most of today’s state-of-the-art generative AI LLM models (e.g., GPT-4 [28] and Llama-2 [35]) consist of a decoder-only Transformer architecture [38]. The decoder-only Transformer models comprise a number of decoder layers stacked on top of each other. Each decoder layer includes a self-attention and a multi-layer perceptron (MLP) layer, each of which are followed by normalization layers. The self-attention layers can further be decomposed into three input projection layers, a positional embedding, a self-attention kernel (e.g., Flash attention [7]), and an output projection layer. Similarly, the MLP layers can be decomposed into three projection layers and one activation layer. A large percentage of memory footprint and computational workload of Transformer architectures comes from the projection layers and self-attention kernels, which puts them at the center of focus for LLM acceleration.
@@
In contrast, the algebraic operations in self-attention layers during the decode phase are mostly matrix-matrix mutliplication with one very narrow matrix (often no wider than 8), thus providing an operational intensity of  16 Op/word while most accelerators’ roofline is at >100 Op/word. Similarly, the data reuse in projection layers is limited to the number of batches in the decode phase, which also often result in low operational intensity. Therefore, in various application scenarios where the number of batches might be limited to a small number due to low rate of incoming queries or limited memory capacity (e.g., mobile platforms or latency-constraint online services), the decode phase suffers from memory-bandwidth bottlenecks. As a result, the autoregressive generative AI applications with LLMs is generally considered a memory-bandwidth bottlenecked process.
@@
To minimize the communication overhead, the self-attention layers are typically partitioned in the attention-head dimension, where the calculation of each attention head is independent of each other up to the output projection layer of the self-attention layers [31]. Similarly, the MLP layers are also partitioned in their intermediate dimension, which corresponds to the columns and rows of the input and output linear projections, respectively, so that much of the calculation can be performed locally. As a result of this partitioning scheme, accelerators need to perform only one allreduce call per self-attention and MLP layer.
@@
Figure 1: Illustration of overlapping weights and KV-cache prefetching with allreduce operations in a decode stage of a Llama architecture [35]. The x-axis shows time in microseconds and reflects the timing of an Ascend 910B device [12].
3 Proposed Method
@@
Figure 1 illustrates the timeline of a decode stage of LLM inference with and without the proposed method. In vanilla execution (top-half of the figure), the query (denoted as $W_{q}$), key, and value linear projections of the attention layers have a data dependency with the preceding allreduce operation, as they need to wait for the reduction of the activations from the previous layer. Similarly, the gate (denoted as $W_{G}$) and up (denoted as $W_{U}$) projections of the MLP layers have also a data dependency with the preceding allreduce operation. As a result, modern ML frameworks (e.g., Pytorch) execute these operations sequentially in a single stream. Unfortunately, the devices remain idle while waiting for allreduce operations, leading to underutilization and long latencies.
@@
While the linear projections need to wait for allreduce operations due to the data dependencies on activations, their weights are read-only, so they do not have any dependencies with the preceding operations. Similarly, the $K$ and $V$ caches in the $QK^{T}$ and $SV$ calculations of the self-attention layers are also read-only in the decode stage, except the last entry that is updated based on the key and value that are calculated in the preceding linear projections. Therefore, the weights and the KV caches can be prefetched before performing the forward pass of these layers.
The bottom-half of the Fig.1 illustrates how the proposed method works. In a parallel stream to the main one (denoted as prefetch stream in Fig.1), devices start prefetching the weights and KV-cache while waiting for the allreduce operations. As a result, the forward pass of the linear projections and the self-attention take shorter, as the weights and KV-cache are hit and quickly read from the L2 cache of these devices. Consequently, the compute resources are better utilized and the execution time of the each decode layer is shortened.
Figure 2: Memory footprint of Attention and MLP layers of various LLMs for varying number of devices. The models are assumed to be distributed using tensor parallelism, and both weights and KV-cache are stored in int8 precision. Batch size and context length are taken as 8 and 16k, respectively. Horizontal lines represent the L2 capacity of various state-of-the-art NPU and GPUs, namely Ascend 910B [12], TPUv5p [36], TPUv5e [37], and H100 [27].
@@
Fig. 2 demonstrates that tensor parallelism up to 16 devices is sufficient to reduce the memory requirements of the Attention and MLP layers of even the largest models down to a level at which they can fit in the L2 cache of the existing commercial AI accelerators. Modern LLM inference systems often consist of many more devices. For instance, Deepspeed-inference scales up to 256 GPUs [3] and Cloud TPUv5e inference instances contain up to 256 chips [37]. Therefore, we conclude that the proposed method is applicable to most of the widely used, open-source LLMs given the scale of the current LLM inference systems and the L2 cache capacity of the commercial AI accelerators.
Figure 3: An overview of the proposed PRESERVE framework.
@@
Many of today’s hardware vendors provide graph optimization frameworks (e.g., CUDA Graphs [25], CANN Graph Engine [15] etc.), which take a program written in a hardware-agnostic high-level programming language (e.g., Pytorch) and perform compiler-level optimizations, such as operator insertion, fusion, removal. Moreover, such graph optimization frameworks typically have the full view of the entire graph, which enables to optimize for best use of memory resources. To that end, we also developed the PRESERVE framework, which inserts prefetching operators for weights and KV-caches into a computational graph of LLM inference, while automatically managing stream synchronization and cache optimizations in order to minimize cache pollution and redundant memory bandwidth usage.
@@
Figure 3 illustrates the overview of the proposed PRESERVE framework. On the host side, the framework takes a user code that implements an LLM model inference written in one of the supported ML frameworks such as Pytorch and Tensorflow. Then, the user code is compiled into an intermediate representation (IR) using either the built-in compiler of the ML frameworks (e.g., TorchDynamo) or one of the hardware-specific compilers (e.g., Ascend ATC [14]). Next, IR is passed to a graph optimizer (e.g., Ascend Graph Engine (GE) [23]), which performs graph-level optimizations and inserts prefetch operators for the suitable compute and communication operations into the graph. Finally, the optimized graph is compiled into an offline model (executable binaries) using a vendor-specific operator library (e.g., Ascend CANN [15]).
@@
Algorithm 1 describes the proposed operator insertion algorithm for weight and KV-cache prefetching. The proposed algorithm takes a computation graph and L2 cache capacity of the target hardware platform as inputs. Then, the algorithm finds all allreduce operations in the computation graph and iterate over each of these allreduce operations. For each allreduce operation, the algorithm visits the children nodes in a breadth-first order until it reaches the end of graph or the next allreduce operation. For each children node of the type MatMul or SelfAttention, the algorithm estimates the required memory size for prefetching and calculates the total memory size of all preceding prefetch operators between the current node and the allreduce operation. If the total memory size is smaller than the L2 cache capacity, a prefetch operator for the weight or KV-cache of the current node is inserted in a parallel stream to the allreduce operation. If the total memory size exceeds the L2 cache capacity, the prefetch operator is not inserted to avoid cache eviction, and the algorithm continues with the next allreduce operation.
@@ 4 Experiments
For the experiments in this section, we choose a number of open-source LLMs that are widely adopted by the community: Llama3-8b, Llama3-70b [35], Qwen2-7B, Qwen2-72B [40], Phi-3-small, and Phi-3-medium [1]. We varied the batch sizes from 1 to 64 and the sequence length from 2k to 32k to explore the impact of batch size and sequence length on the performance of the proposed method. For sake of simplicity, we assume static batching with equal sequence lengths. The prefill and decode lengths are taken as the 2/3 and 1/3 of the sequence length, respectively.
All benchmarks are implemented in Pytorch with the torch-npu backend [13], based on the reference implementation given in torchair library [16] and compiled using TorchDynamo. The parallel execution is achieved via the communication primitives provided by the torch.distributed library. The proposed method is implemented using the Graph Engine (GE) feature of the Compute Architecture for Neural Networks (CANN) [15], which enables to perform weight and KV-cache prefetching during inference.
The experiments are performed on a Huawei Atlas 800T A2 server with 4x Kunpeng 920 48-core CPUs and 8x Ascend 910B NPUs. Each NPU has an L2 cache capacity of 192 MB and is equipped with 24 DaVinci AI cores [34] and 64 GB HBM memory, which provides a total theoretical FLOPs of 800 TOPS/s in int8 precision and 1.6TB/s off-chip memory bandwidth [12, 21]. The Ascend 910B NPUs in the server are tightly interconnected to each other through an HCCS interconnect fabric in a full-mesh topology.
Table 1: End-to-end inference execution time in seconds with the baseline and proposed methods for various benchmarks and number of NPUs. Batch size and maximum sequence length are taken as 4 and 16k, respectively. Both activations and weights are in int8 format.
@@
Figure 4: Speedups obtained with the proposed method for varying batch size and maximum sequence length for various models. Both activation and weights are in int8 precision. The experiments are performed on four NPUs. N/A indicates insufficient HBM capacity.
@@
In conclusion, in this section, we evaluated the effectiveness and the performance of the proposed method for various LLMs under diverse inference settings, such as the number of NPUs, batch size, and sequence lengths. Our experiments demonstrated that the proposed method achieves up to 1.61$\times$ end-to-end speedup on commercial AI accelerators. Our study on batch size and sequence length identified the optimal settings where the proposed method is most effective and showed that the proposed method achieves significant speedups under a wide range of inference settings.
5 Design Space Exploration
The experiment results in the previous section demonstrate significant improvement in the performance of distributed LLM inference with PRESERVE framework on existing commercial AI accelerators. However, various hardware design parameters and specifications have an impact on the performance of the PRESERVE framework. Therefore, in this section, we investigate how the hardware design parameters (i.e., L2 cache size, L2 bus bandwidth, cluster size, network bandwidth) affect the effectiveness of the PRESERVE framework and perform a design space exploration to identify the optimal hardware design parameters that maximize the inference performance of AI accelerators.

### v2

@@
Abstract.
Large language models (LLMs) are typically served from clusters of GPUs/NPUs that consist of large number of devices. Unfortunately, communication between these devices incurs significant overhead, increasing the inference latency and cost while limiting the scalability. Prior work addressed this issue by overlapping communication with compute, but has severe limitations due to the data dependencies between these operations. In this paper, we propose PRESERVE, a novel framework that prefetches model weights and KV-cache from off-chip HBM memory to the on-chip cache of AI accelerators during the communication operations, which offers various advantages and performance improvements compared to prior methods.
Through extensive experiments conducted on commercial AI accelerators, we demonstrate up to 1.6× end-to-end speedup on state-of-the-art, open-source LLMs. Additionally, we perform a design space exploration that identifies the optimal hardware configuration for the proposed method, showing a further 1.25× improvement in performance per cost by selecting the optimal L2 cache size. Our results show that PRESERVE has the potential to mitigate the memory bottlenecks and communication overheads, offering a solution to improve the performance and scalability of the LLM inference systems.
LLM inference, distributed systems, compute-communication overlap, L2 prefetching
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: EuroSys: European Conference on Computer Systems; April 13–16, 2026; Edinburgh, Scotlandsubmissionid: 183ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Large language models (LLM) have been widely deployed across various services and application domains such as chat assistants (OpenAI, 2023), code generation (Chen et al., 2021), and knowledge retrieval (Lewis et al., 2020). While LLMs have demonstrated outstanding capabilities in these domains, their substantial computational requirements result in slow and costly inference. To address these challenges, specialized accelerators have been developed and deployed for LLM inference. Consequently, the performance, efficiency, and scalability of LLM inference accelerators are critical to the speed and cost of the LLM inference.
Figure 1. Conceptual comparison between A) vanilla inference, B) inference with GEMM+Allreduce fused kernels (Rashidi et al., 2021; Hoefler and Lumsdaine, 2008; Punniyamurthy et al., 2024; Chang et al., 2024; Wang et al., 2023), C) PRESERVE (this work).
@@ 1. Introduction
To mitigate the communication overhead, prior studies have proposed fusing matmul and allreduce operations to overlap compute and communication (Rashidi et al., 2021; Hoefler and Lumsdaine, 2008; Punniyamurthy et al., 2024; Chang et al., 2024; Wang et al., 2023), as shown in Figure 1. While these methods might improve the resource utilization by overlapping one GEMM and one consecutive allreduce operations, their effectiveness is limited as they can only fuse two consecutive operations, which is often insufficient to hide the entire communication latency in large inference clusters. Moreover, these methods can not be applied to KV cache, because there are many other operations between self-attention and allreduce operations in a typical LLM architecture. Finally, operator fusion requires high engineering efforts and expertise in low-level kernel development, which also reduces the practicality of these methods. As a result, existing compute-communication overlap methods fail to provide an efficient, effective and comprehensive solution.
In this paper, we address the communication overhead in distributed LLM inference by exploiting the ever-growing on-chip memory capacity in modern AI accelerators (e.g., NVIDIA GB200: 126 MB of L2 cache (NVIDIA, 2025), AMD MI300X: 256 MB of L3 cache (AMD, 2025), Huawei Ascend 910B: 196 MB of L2 cache (Huawei, 2022)). We propose prefetching the model weights and KV-cache from off-chip HBM memory to on-chip cache of the AI accelerators during the communication operations to overlap memory reads for the model weights and KV-cache with communication, hiding the latency of the latter. Once the communication operations are completed, compute cores access the required data from the on-chip cache, no longer limited by off-chip memory bandwidth. As a result, the overall decoding phase of LLM inference is executed faster and more efficiently, reducing end-to-end latency and execution cost.
Due to the data dependencies between the communication operations and the subsequent matmul operations, neither hardware schedulers nor modern ML compiler can automatically and trivially perform prefetching on model weights and KV-cache. Moreover, if not handled carefully, prefetching excessive amounts of data poses the risk of polluting the cache, which may result in a slowdown rather than a speedup. Therefore, we propose PRESERVE, a novel method that prefetches model weights and KV-cache from off-chip HBM memory to the on-chip device cache in parallel to the collective communication operations. To maximize the performance gains without the need for any modification in the user code for ease of development, we implemented PRESERVE as a graph optimization framework. Our graph optimization scheme inserts prefetch instructions into a given computation graph, to be executed in parallel streams to the collective communication operations. PRESERVE also keeps track of the prefetched data and estimates L2 cache usage during compilation time in a way that the total prefetched data size at any time does not exceed a predefined upper limit in order to prevent cache pollution and potential slowdowns.
@@ 2. Background
Most of today’s state-of-the-art generative AI LLM models (e.g., GPT-4 (OpenAI, 2023) and Llama-2 (Touvron et al., 2023)) consist of a decoder-only Transformer architecture (Vaswani et al., 2017). The decoder-only Transformer models comprise a number of decoder layers stacked on top of each other. Each decoder layer includes a self-attention and a multi-layer perceptron (MLP) layer, each of which are followed by normalization layers. The self-attention layers can further be decomposed into three input projection layers, a positional embedding, a self-attention kernel (e.g., Flash attention (Dao et al., 2022)), and an output projection layer. Similarly, the MLP layers can be decomposed into three projection layers and one activation layer. A large percentage of memory footprint and computational workload of Transformer architectures comes from the projection layers and self-attention kernels, which puts them at the center of focus for LLM acceleration.
@@
In contrast, the algebraic operations in self-attention layers during the decode phase are mostly matrix-matrix mutliplication with one very narrow matrix (often no wider than 8), thus providing an operational intensity of  16 Op/word while most accelerators’ roofline is at ¿100 Op/word. Similarly, the data reuse in projection layers is limited to the number of batches in the decode phase, which also often result in low operational intensity. Therefore, in various application scenarios where the number of batches might be limited to a small number due to low rate of incoming queries or limited memory capacity (e.g., mobile platforms or latency-constraint online services), the decode phase suffers from memory-bandwidth bottlenecks. As a result, the autoregressive generative AI applications with LLMs is generally considered a memory-bandwidth bottlenecked process.
@@
Figure 2. Memory footprint of Attention and MLP layers of various LLMs for varying number of devices. The models are assumed to be distributed using tensor parallelism, and both weights and KV-cache are stored in int8 precision. Batch size and context length are taken as 8 and 16k, respectively. Horizontal lines represent the L2 capacity of various state-of-the-art NPU and GPUs, namely Ascend 910B (Huawei, 2022), TPUv5p (Vahdat and Lohmeyer, 2023a), TPUv5e (Vahdat and Lohmeyer, 2023b), and H100 (NVIDIA, 2023b).
@@
To minimize the communication overhead, the self-attention layers are typically partitioned in the attention-head dimension, where the calculation of each attention head is independent of each other up to the output projection layer of the self-attention layers (Shoeybi et al., 2019). Similarly, the MLP layers are also partitioned in their intermediate dimension, which corresponds to the columns and rows of the input and output linear projections, respectively, so that much of the calculation can be performed locally. As a result of this partitioning scheme, accelerators need to perform only one allreduce call per self-attention and MLP layer.
@@
3. Proposed Method
@@
In vanilla execution, the query, key, and value linear projections of the attention layers have a data dependency with the preceding allreduce operation, as they need to wait for the reduction of the activations from the previous layer. Similarly, the gate and up projections of the MLP layers also have a data dependency on the preceding allreduce operation. As a result, modern ML frameworks (e.g., Pytorch) execute these operations sequentially in a single stream. Unfortunately, the devices remain idle while waiting for allreduce operations, leading to underutilization and long latencies.
@@
While the linear projections need to wait for allreduce operations due to data dependencies on activations, their weights are read-only, so they do not have any dependencies with the preceding operations. Similarly, the $K$ and $V$ caches in the $QK^{T}$ and $SV$ calculations of the self-attention layers are also read-only in the decode stage, except for the last entry, which is updated based on the key and value calculated in the preceding linear projections. Therefore, the weights and the KV caches can be prefetched before performing the forward pass of these layers.
With the proposed method, devices start prefetching the weights and KV-cache in a parallel stream from off-chip HBM memory to on-chip buffers (e.g., L2 cache) while waiting for the allreduce operations. When the allreduce operation is completed and the device is ready to execute the next layers, the weights and KV-cache are already fetched to the L2 cache, which typically provides 5-10$\times$ higher bandwidth than off-chip HBM memories. As a result, the forward pass of the linear projections and the self-attention takes much shorter, improving resource utilization and speeding up the decode phase.
@@
Fig. 2 demonstrates that tensor parallelism up to 16 devices is sufficient to reduce the memory requirements of the Attention and MLP layers of even the largest models down to a level at which they can fit in the L2 cache of the existing commercial AI accelerators. Modern LLM inference systems often consist of many more devices. For instance, Deepspeed-inference scales up to 256 GPUs (Aminabadi et al., 2022) and Cloud TPUv5e inference instances contain up to 256 chips (Vahdat and Lohmeyer, 2023b). Therefore, we conclude that the proposed method is applicable to most of the widely used, open-source LLMs given the scale of the current LLM inference systems and the L2 cache capacity of the commercial AI accelerators.
Figure 3. An overview of the proposed PRESERVE framework.
@@
Many of today’s hardware vendors provide graph optimization frameworks (e.g., CUDA Graphs (NVIDIA, 2019), CANN Graph Engine (Huawei, 2024c) etc.), which take a program written in a hardware-agnostic high-level programming language (e.g., Pytorch) and perform compiler-level optimizations, such as operator insertion, fusion, removal. Moreover, such graph optimization frameworks typically have the full view of the entire graph, which enables to optimize for best use of memory resources. To that end, we also developed the PRESERVE framework, which inserts prefetching operators for weights and KV-caches into a computational graph of LLM inference, while automatically managing stream synchronization and cache optimizations in order to minimize cache pollution and redundant memory bandwidth usage.
@@
Figure 3 illustrates the overview of the proposed PRESERVE framework. On the host side, the framework takes a user code that implements an LLM model inference written in one of the supported ML frameworks such as Pytorch and Tensorflow. Then, the user code is compiled into an intermediate representation (IR) using either the built-in compiler of the ML frameworks (e.g., TorchDynamo) or one of the hardware-specific compilers (e.g., Ascend ATC (Huawei, 2024b)). Next, IR is passed to a graph optimizer (e.g., Ascend Graph Engine (GE) (Liao et al., 2019)), which performs graph-level optimizations and inserts prefetch operators for the suitable compute and communication operations into the graph. Finally, the optimized graph is compiled into an offline model (executable binaries) using a vendor-specific operator library (e.g., Ascend CANN (Huawei, 2024c)).
@@
Algorithm 1 describes the proposed operator insertion algorithm for weight and KV-cache prefetching. The proposed algorithm takes a computation graph and L2 cache capacity of the target hardware platform as inputs. Then, the algorithm finds and iterates over all communication operations in the computation graph. For each communication operation, the algorithm visits the children nodes in a breadth-first order until it reaches the end of graph or the next communication operation. For each children node of the type MatMul or SelfAttention, the algorithm estimates the required memory size for prefetching and calculates the total memory size of all preceding prefetch operators between the current node and the communication operation. If the total memory size is smaller than the L2 cache capacity, a prefetch operator for the weight or KV-cache of the current node is inserted in a parallel stream to the communication operation. If the total memory size exceeds the L2 cache capacity, the prefetch operator is not inserted to avoid cache eviction, and the algorithm continues with the next communication operation.
@@ 4. Experiments
For the experiments in this section, we choose a number of open-source LLMs that are widely adopted by the community: Llama3-8b, Llama3-70b (Touvron et al., 2023), Qwen2-7B, Qwen2-72B (Yang et al., 2024), Phi-3-small, and Phi-3-medium (Abdin et al., 2024). We varied the batch sizes from 1 to 64 and the sequence length from 2k to 32k to explore the impact of batch size and sequence length on the performance of the proposed method. For sake of simplicity, we assume static batching with equal sequence lengths. The prefill and decode lengths are taken as the 2/3 and 1/3 of the sequence length, respectively.
All benchmarks are implemented in Pytorch with the torch-npu backend (Huawei, 2024a), based on the reference implementation given in torchair library (Huawei, 2024d) and compiled using TorchDynamo. The parallel execution is achieved via the communication primitives provided by the torch.distributed library. The proposed method is implemented using the Graph Engine (GE) feature of the Compute Architecture for Neural Networks (CANN) (Huawei, 2024c), which enables to perform weight and KV-cache prefetching during inference.
The experiments are performed on a Huawei Atlas 800T A2 server with 4x Kunpeng 920 48-core CPUs and 8x Ascend 910B NPUs. Each NPU has an L2 cache capacity of 192 MB and is equipped with 24 DaVinci AI cores (Tang and Wang, 2023) and 64 GB HBM memory, which provides a total theoretical FLOPs of 800 TOPS/s in int8 precision and 1.6TB/s off-chip memory bandwidth (Huawei, 2022; Li et al., 2024). The Ascend 910B NPUs in the server are tightly interconnected to each other through an HCCS interconnect fabric in a full-mesh topology.
Table 1. End-to-end inference execution time in seconds with the baseline and proposed methods for various benchmarks and number of NPUs. Batch size and maximum sequence length are taken as 4 and 16k, respectively. Both activations and weights are in int8 format.
@@
Figure 4. Speedups obtained with the proposed method for varying batch size and maximum sequence length for various models. Both activation and weights are in int8 precision. The experiments are performed on four NPUs. N/A indicates insufficient HBM capacity.
@@
Comparison to compute-communication overlap fused kernels
To demonstrate the performance benefits of the proposed method over the prior methods using fused kernels that overlap GEMM and Allreduce operations, we now compare the performance of PRESERVE over a baseline, in which the last linear layer of each Attention and MLP layer, namely $W_{out}$ and $W_{down}$ respectively, are fused with the succeeding Allreduce operations. To fuse the GEMM and Allreduce operations, we use the npu_mm_all_reduce_base (Huawei, 2025) operator from the torch_npu library. We then measure the execution time of Llama3-70B (fp16) decoding with both PRESERVE and the baseline, on a cluster of 4x Ascend 910B NPUs for batch sizes ranging from 16 to 1024 and maximum sequence lengths ranging from 128 to 8192, both incrementing as powers of two.
Figure 5 shows that PRESERVE outperforms the baseline at small batch sizes while the baseline performs better at large batch sizes. There are multiple reasons why PRESERVE is advantageous over the fused Matmul+Allreduce operators at small batch sizes. To begin with, PRESERVE allows overlapping the HBM reads of any weights and KV-cache with communication, whereas the Matmul+Allreduce operators can only fuse consecutive operations, which limits their applicability only to the last projection layers of each layer, omitting the KV-cache entirely. Moreover, the fused operators typically partition the matmul operations into multiple chunks and pipeline the execution to hide the latency of the communication. However, small chunks of data often lead to inefficiencies in memory accesses, which inhibit the performance of these operators at small batch sizes. As a result, PRESERVE outperforms the baseline with fused Matmul+Allreduce operators at small batch sizes with a speedup up to 1.19×1.19$\times$1.19 ×.
In contrast, due to higher computational intensity and the growing size of KV-cache (as already discussed before and shown in Figure 4), the effectiveness of PRESERVE slowly diminishes with the increasing batch size. As a result, the baseline becomes faster than PRESERVE at batch sizes equal to or larger than 512, shown by the less-than-one speedups in Figure 5, with a minimum speedup of 0.93×0.93$\times$0.93 × at the batch size of 1024 and max. sequence length of 128. Therefore, we conclude that PRESERVE is the better choice in scenarios, in which the batch size is smaller than 512 (e.g., online and resource-constraint inference), whereas the fused Matmul+Allreduce operators proposed by the prior work remain as the method of choice for scenarios that require batch sizes equal to or larger than 512 (e.g., offline inference).
Figure 5. Comparison over the fused GEMM+Allreduce operator baseline for various batch size (y-axis) and max. sequence lengths (x-axis) for Llama3-70b in fp16. The red-colored squares indicate that PRESERVE is faster than the fused GEMM+Allreduce. N/A denotes out-of-memory.
In conclusion, in this section, we evaluated the effectiveness and the performance of the proposed method for various LLMs under diverse inference settings, such as the number of NPUs, batch size, and sequence lengths. Our experiments demonstrated that the proposed method achieves up to 1.61$\times$ end-to-end speedup on commercial AI accelerators. Our study on batch size and sequence length identified the optimal settings where the proposed method is most effective and showed that the proposed method achieves significant speedups, outperforming the baseline methods under a wide range of inference settings.
5. Design Space Exploration
The experiment results in the previous section demonstrate significant improvement in the performance of distributed LLM inference with PRESERVE framework on existing commercial AI accelerators. However, various hardware design parameters and specifications have an impact on the performance of the PRESERVE framework. Therefore, in this section, we investigate how the hardware design parameters affect the effectiveness of the PRESERVE framework and perform a design space exploration to identify the optimal hardware design parameters that maximize the inference performance of AI accelerators.